OpenStudy (ookawaiioo):

There is 3,000,000,000 base pairs and it needs to be replicated. The enzyme can replicate 60 base pairs/sec. If replication is finished in 20 mins (1200sec), how many origin of replications are there?

OpenStudy (hoblos):

each origin replicates 60 * 1200 = 72000 base pairs then we need 3000000000/72000=41667 origins of replication

OpenStudy (ookawaiioo):

THANKS! I got the same answer! You a good man

OpenStudy (anonymous):

A = Pert for t: 60=e^1200sec*k 60=2.718*k k=22.075 for the origin: 3,000,000,000=Pe1200*22.075 3,000,000,000=Pe^26490 3,000,000,000=P(2.718) P=

OpenStudy (hoblos):

welcome